OpenStudy (anonymous):

Gabriel is completing a science fair project on water conservation by filling a bucket with water from one of his house’s downspouts. During a rainstorm, at a flow rate of 0.05 liter per second, the bucket is filled in 100 seconds. During the next rain, the flow rate is 0.2 liter per second, and the bucket is filled in 25 seconds. Which equation represents the relationship between the flow rate, f, and the time, t, it takes to fill the bucket?

OpenStudy (anonymous):

[A] t = 2000f [B] t = 5f [C] t= 5/f [D] t = f/2000